OpenStudy (anonymous):

Solve the equation. 2(x-2)^2 +x^2 = x(x+50)-48x

OpenStudy (anonymous):

I thought it was no solution, but the answer is apparently that the solution set is 1 and 4, but I don't see how

OpenStudy (anonymous):

You want the answer or steps?

OpenStudy (anonymous):

The steps

OpenStudy (anonymous):

\[2(x-2)^2 + x^2 = x(x+50) - 48x\] \[2( x^2 -4x +4) + x^2 = x^2 + 50x - 48x\]

OpenStudy (anonymous):

get that?

OpenStudy (anonymous):

Not really. Why did you rewrite them that way?

OpenStudy (anonymous):

cause i have to, (x-2)^2 = (x-2)(x-2) and that makes the x^2 - 4x + 4

OpenStudy (anonymous):

Oh right

OpenStudy (anonymous):

and x(x + 50) = x*x + x*50 = x^2 + 50x

OpenStudy (anonymous):

right

OpenStudy (anonymous):

\[2x^2 -8x +8 + x^2 = x^2 +2x\] \[3x^2 - 8x + 8 -x^2 -2x=0\]

OpenStudy (anonymous):

how bout that?

OpenStudy (anonymous):

\[2x^2 -10x +8=0\] \[2(x^2-5x+4)=0\]

OpenStudy (anonymous):

Alright, yeah, I got that part with the left side

OpenStudy (anonymous):

2(x-4)(x-1)=0 x = 4 OR x = 1

OpenStudy (anonymous):

Oh alright. Thanks. There wasn't anything of the sort in my notes but I remember doing something like it in class.

OpenStudy (anonymous):

ghe, enjoy :)